# Generating Jackiw-Teitelboim Euclidean gravity from static three-dimensional Maxwell-Chern-Simons electromagnetism

### v1

@@ 3 Achieving gravity and the flat solution
In this section we map the static electromagnetic action (2.16) into a two-dimensional gravity theory in manifolds of Euclidean signature. The details and formal aspects of the mapping can be found in [55, 56]. In those papers, extra auxiliary non-physical fields are introduced in order to make the mapping well defined. In here, we omit such particularities because the same method applies in the present case. In a nutshell, the mapping is based on the isomorphism between the gauge group $U(1)$ of electrodynamics and the $SO(2)$ rotations of Euclidean two-dimensional gravity. The map between the fields of electrodynamics and gravity is [^1] The factor $e^{4}$ is different from the pure two-dimensional case [55, 56] because the canonical dimension of the electric charge is different. In fact, if $D$ is spacetime dimension, the canonical dimension of the electric charge is $[e]=(4-D)/2$.,
[…]
with $e^{a}=e^{a}_{i}dX^{i}$ being the zweibein 1-form, with vanishing mass dimension $[e]=0$, while $\omega^{ab}=\omega^{ab}_{i}dX^{i}$ is the spin-connection, with mas dimension given by $[\omega]=1$. The point $x\in\mathbb{R}^{2}$ is a point in the Euclidean space where action (2.16) is defined and $X\in M^{2}$ is a point in the two-dimensional manifold $M^{2}$ where gravity will emerge. The indices $a,b,c,\ldots,h\in\{1,2\}$ are associated with the tangent space of $M^{2}$, essentially a local inertial frame, while $i,j,k,\ldots\in\{1,2\}$ refers to a general coordinate system of the manifold $M^{2}$. The essence of the mapping is the identification of the fields of three-dimensional static electromagnetism with the geometric fields of the manifold, i.e., the dynamics of the electromagnetic fields are absorbed by the manifold.
Performing the mapping (3.1) in the action (2.16), one straightforwardly finds
$S_{2grav}=\int\left\{-\frac{1}{e^{2}}\left(d\Phi\ast d\Phi+2e^{4}\epsilon_{ab}R^{ab}-e^{8}\epsilon_{ab}e^{a}e^{b}\right)+\kappa\left[2\Phi\epsilon_{ab}R^{ab}-\epsilon_{ab}d\left(\Phi\omega^{ab}\right)\right]\right\}\;,$ (3.2)
with $R^{ab}=d\omega^{ab}$ being the curvature 2-form and $\epsilon_{ab}$ being the Levi-Civita symbol. Therefore, inferring the parameter identifications [^2] Obviously, $[G]=-1$ and $[\Lambda]=1$.,
$G$ $=$ $\frac{1}{16\pi e^{2}}\;,$
$\Lambda$ $=$ $e^{2}\;,$ (3.3)
@@
This is a two-dimensional gravity action, in the first-order formalism, coupled with a scalar field, naturally identified as the dilaton. The constants $G$ and $\Lambda$ are recognized as the two-dimensional Newtonian and cosmological constants, respectively. Remarkably, since we know the value of the electric charge, we can compute the numerical values of the gravitational constants through expressions (3.3). Indeed,
@@
The first term of the action (3.4) is the standard kinematic term for a real scalar field, up to a constant factor. The second and third pieces are the typical Einstein-Hilbert (EH) and cosmological constant terms. The EH action, in two dimensions, is known to be of topological nature, a property evidenced by the fact that it is a total derivative. In fact, in two dimensions, the EH term coincides with the GB term. See Section 5.
Continuing the analysis of the action (3.4), the next term is the dilaton-curvature coupling, emerging from the CS term. The last term is just a total derivative. This term, just like the EH one, could only contribute to the field equations at the boundary if this boundary is non-trivial.
Clearly, the action (3.4) is, up to boundary terms, manifestly gauge invariant under $SO(2)$ rotations in tangent space,
@@
Thus, the action (3.4) is a genuine Euclidean two-dimensional gravity, in the FOF, coupled to a scalar field. We point out that transformations (3.7) are directly obtained from (2.15), (3.1) and the identification $\alpha(x)=\epsilon_{ab}\alpha^{ab}(X)$, see [55, 56]. We also stress that action (3.4) displays diffeomorphism symmetry, which is implicit in form notation.
In the next section, we will show that this action can be transmuted in JT-type gravities. For now, let us just exploit the dynamics of the action (3.4). The field equation of $\Phi$ is just
@@
suggesting that the scalar field carries some dynamics. It also suggests that the scalar field affects non-trivially the manifold geometry. However, the spin-connection equation gives
@@
So, the dilaton solution is simply an arbitrary constant field configuration, possibly zero. Furthermore, combining equations (3.8) and (3.9) one easily finds that the manifold is locally flat,
@@
It remains to derive the equation of the zweibein. The only non-trivial term remaining in the action (3.4) that depends on the zweibein is the area term. Then, the final equation is simply the minimization of the area of the manifold (Plateau problem), just like two-dimensional Mardones-Zanelli gravity [33, 34]. The solution here is thus a flat space with minimal area. Of course, the specific manifold depends on the boundary conditions. Typical non-compact examples are planes, cylinders, and disks. In the case of compact manifolds, the flat torus is the standard solution. See Figure 1, in Section 5, for examples. In fact, we will return to the discussion of these surfaces in Section 5.
@@ 4.1 Hyperbolic solution
The first way to obtain a JT gravity is by carrying out the curvature shift
@@
enforcing the JT usual term to directly appear in the gravity action (3.4). However, before we perform this shift in the proper way, let us define it a bit more rigorously. In fact, the shift (4.1) is actually generated by a shift in the spin-connection, namely
@@
meaning that the field $X$ must be a non-conservative field with a canonical dimension of $[X]=-1$. For further reference, we name condition (4.3) shift constraint. Thus, once the zweibein $e$ describing the manifold is chosen (note that there are infinite possibilities due to the $SO(2)$ gauge symmetry and diffeomorphisms), the field $X$ can be determined. Integration of equation (4.3), and resorting to Stokes’ theorem, gives the global condition
@@
Another point of importance concerning equation (4.3) is that it is not gauge covariant unless $dX$ is covariant, which is actually not true. In fact, the shift (4.2) suggests that $X$ is a gauge invariant field. As a consequence, gauge symmetry is broken by the constraint (4.3); it may be seen, thus, as a gauge fixing for the model. In fact, together with the shift (4.2), it can be incorporated into the action (3.4) with a suitable 0-form Lagrange multiplier [^3] The Lagrange multiplier carry mass dimension 3, $[\lambda]=3$. $\lambda_{ab}$ by means of the addition of the action
@@
This term clearly breaks gauge symmetry, assuming that $\lambda$ is either gauge invariant or covariant.
@@
Let us derive the field equations. The equation for the Lagrange multiplier naturally imposes the shift constraint (4.3). The equation of the spin-connection remains the same, equation (3.9). Thus, $\Phi$ is again undetermined. It is constant, but not necessarily zero. However, nothing forbids us to choose $\Phi=0$. The equation for $X$, provides (already employing (3.9))
@@
which is just JT solution for negative constant curvature in Euclidean spaces.
@@
This is actually quite an interesting result because a Lagrange multiplier works as a force obliging the system to obey the constraint. It turns out that the cosmological constant is that force. In essence, the cosmological constant is the force that bends the manifold.
The solution we found so far is just a dilatonic trivial vacuum and a manifold with negative constant curvature, i.e., a hyperbolic surface. In two-dimensional Euclidean spaces, such manifolds, if compact, are the connected $n$-holed torii [63, 64] - See Figure 3 and Section 5 for further discussions.
@@ 4.2 Spherical solution
Let us compute the field equations. Equation (3.9) remains valid, stating that $\Phi$ is a constant field again. The equation of $\lambda$ simply enforces the shift constraint (4.3). Again, the equation of the field $X$ gives that $d\lambda=0$, due to (3.9). Thus, $\lambda$ is also a constant field. The field equation of the dilaton, already employing the equations of $\omega$ and $\lambda$, gives the JT solution for a spherical manifold (constant positive curvature),
@@ 4.3 Extra remarks about the shift constraint
To finalize this Section, we remark that the shift constraint (4.3) can also be visualized as the imposition of a background. Therefore, the gauge fixing is actually a choice of an appropriate background. The only difference between the hyperbolic and the spherical cases is just the relative sign of the background. Essentially, the shifts (4.2) and (4.12), together with the constraint (4.3), impose the background to be hyperbolic or spherical. Moreover, since in two dimensions the EH term is topological, there is no room to generate Einstein-like equations in order to provide dynamics around the background.
In fact, the shifts (4.2) and (4.12) are simply stating that ±$\sim\pm X$ is the spin-connection background in each case. The shift constraint (4.3) states that $X$ must be the field that generates the specific (hyperbolic or spherical) background curvature. Moreover, it is known that the spin connection carries a specific term depending on the zweibein. Therefore, it is only natural that the shift constraint is written in terms of the zweibein. This is even evidenced by the fact that torsion is zero in our solutions.
@@
It is also interesting to notice that the JT term could emerge directly from the mapping by considering an electromagnetic background field. Right from the action (2.16), one could consider the shift $F\longrightarrow F\pm\Theta/2$. Or, equivalently, a shift on the electromagnetic field $a\longrightarrow a\pm M$ with $M$ being fixed by $dM=\Theta/2$. In this way, the target gravity action would already contain the usual JT term $\Phi(R\pm\Lambda^{2}/2)$. Therefore, the shift we proposed is equivalent to the existence of a background electromagnetic field in the original MCS theory.
Another comment is that the shift constraint (4.3) induces a relation between the spin-connection and the field $X$. This relation is actually generated by the cohomology of the nilpotent exterior derivative $d$, the well-known de Rham’s cohomology [65]. In fact, the constraint (4.3) and the curvature equations (4.8) (negative sign) and (4.14) (positive sign) can be combined to deduce the on-shell relation
@@
Moreover, since $X$ is gauge invariant, $Y$ must transform as a connection in order to preserve the gauge structure of relation (4.16), namely
@@
Therefore, the gauge invariant field $X$ is essentially a difference between two connections, $X=\mp 2(\omega-Y)/\Lambda^{2}$, i.e., the difference between the spin-connection and an arbitrary closed 1-form.
Condition (4.17) is actually a traditional cohomology problem whose solution is simply given by its non-trivial and trivial parts in de Rham’s cohomology, namely
@@
while $N$ is an arbitrary 0-form (consequently, it cannot be written as an exact quantity) with dimension $[N]=0$. The gauge transformations of $Z$ and $N$ can be inferred from (4.18) and (4.19). Clearly, $Z$ can be chosen to be a gauge connection while $N$ is set to be gauge invariant [^4] Alternatively, there is also the possibility to choose $Z$ to be a gauge invariant field while $N$ would transform in a non usual way by $\delta N=\alpha$, i.e., $N$ would translate under $SO(2)$ gauge transformations. Such transformation indicates that $N$ could be interpreted as a field of angular nature. In this case, the relation (4.19) could be seen as a gauge transformation of the field $Y$ while $N$ would be the gauge parameter.,
@@
The determination of the spin-connection works as follows. Once the zweibein describing the manifold is chosen, one can find any $X$ satisfying the shift constraint (4.3), a closed non-exact field $Z$, satisfying (4.20), and any gauge invariant 0-form $N$. Then, from relations (4.16) and (4.19), the spin-connection is completely determined. The same reasoning works to find $X$ by solving equation (4.8) (or (4.14)) for the spin-connection. In essence, the spin-connection and the field $X$ differ only by any non-trivial field $Z$ plus an exact field. Moreover, $Z$ and $N$ take no part in the solutions we found for JT gravity, see Equations (4.8) and (4.14). Alternatively, $X$ and $\omega$ can be deduced independently from equations (4.3) and (4.8) (or (4.14)), respectively. Then, $Z$ and $N$ can be inferred by comparing $X$ and $\omega$ via relation (4.16).
It is also important to stress that $Y$ must be non-vanishing because of the gauge transformations of $\omega$, $Y$ (see (3.7) and (4.17)), and the gauge invariant field $X$. Otherwise, relation (4.16) would not be gauge covariant. Thus, if $Z\neq 0$, then $dN$ can be set to vanish. On the other hand, since $N$ is gauge invariant, $Z$ cannot be set to zero. Therefore, from now on, we are allowed to fix $N=0$, with no loss of generality. As an alternative, if we do not care about gauge covariance (which is a valid choice since the presence of $X$ in the action already breaks gauge symmetry), we can set $Z=N=0$ and work with a gauge invariant spin-connection. However, we will not follow that path.
@@
Fields $e^{a}$ $\omega^{ab}$ $\Phi$ $X$ $Y$ $Z$ $N$ $\lambda$ $G$ $\Lambda$ $\kappa$
Dimension 00 $1$ $1$ $-1$ $1$ $1$ 00 $3$ $-1$ $1$ 00
Rank $1$ $1$ 00 $1$ $1$ $1$ 00 00 00 00 00
@@
The set of fundamental fields living in $M^{2}$ is thus =$\mathcal{F}=\{e,\omega,X,Z,\Phi,\lambda\}$, interlinked by the field equations
@@
$R^{ab}$ $=$ ±$\pm\frac{\Lambda^{2}}{2}e^{a}e^{b}\;,$
$\omega^{ab}$ $=$ ±$\pm\frac{\Lambda^{2}}{2}X^{ab}+Z^{ab}\;.$ (4.22)
The third equation in (4.22) establishes that, among $\omega$, $X$, and $Z$, we can neglect one of them in favor of the other two. The space of fundamental fields is then reduced, for instance, to $\mathcal{F}_{o}=\{e,X,Z,\Phi,\lambda\}$, where we have chosen to omit the spin-connection.

### v2

@@ 3 Achieving gravity and the flat solution
In this section, we map the static electromagnetic action (2.16) onto a two-dimensional gravity theory defined on manifolds with Euclidean signature. The technical details and formal aspects of the mapping can be found in [56, 57]. In those works, extra auxiliary non-physical fields are introduced in order to make the mapping well defined. Here, we omit such particularities because the same method applies in the present case. In short, the mapping is based on the isomorphism between the gauge group $U(1)$ of electrodynamics and the $SO(2)$ rotations of two-dimensional Euclidean gravity. The map between the fields of electrodynamics and gravity is given by [^1] The factor $\mathrm{e}^{4}$ is different from the pure two-dimensional case [56, 57] because the canonical dimension of the electric charge is different. In fact, if $D$ is spacetime dimension, the canonical dimension of the electric charge is $\mathrm{e}]=(4-D)/2$.
$\Theta(x)$ $\longmapsto$ $\mathrm{e}^{4}\epsilon_{ab}e^{a}(X)e^{b}(X)\;,$
$A(x)$ $\longmapsto$ $\epsilon_{ab}\omega^{ab}(X)\;,$
$\Phi(x)$ $\longmapsto$ $\Phi(X)\;,$ (3.1)
with $e^{a}=e^{a}_{i}dX^{i}$ being the zweibein 1-form, with vanishing mass dimension $[e]=0$, while $\omega^{ab}=\omega^{ab}_{i}dX^{i}$ is the spin connection, with mas dimension given by $\omega]=1$. The point $x\in\mathbb{R}^{2}$ belongs to the Euclidean space where action (2.16) is defined and $X\in M^{2}$ is a point in the two-dimensional manifold $M^{2}$ where gravity will emerge. The indices $a,b,c,\ldots,h\in\{1,2\}$ are associated with the tangent space of $M^{2}$, essentially a local inertial frame, while $i,j,k,\ldots\in\{1,2\}$ refer to a general coordinate system of the manifold $M^{2}$. The core idea of the mapping is the identification of the fields of three-dimensional static electromagnetism with the geometric fields of the manifold; in other words, the dynamics of the electromagnetic fields are absorbed into the manifold.
Performing the mapping (3.1) in the action (2.16), one straightforwardly obtains
$S_{2grav}=\int\left\{-\frac{1}{\mathrm{e}^{2}}\left(d\Phi\ast d\Phi+2\mathrm{e}^{4}\epsilon_{ab}R^{ab}-\mathrm{e}^{8}\epsilon_{ab}e^{a}e^{b}\right)+\kappa\left[2\Phi\epsilon_{ab}R^{ab}-\epsilon_{ab}d\left(\Phi\omega^{ab}\right)\right]\right\}\;,$ (3.2)
where $R^{ab}=d\omega^{ab}$ is the curvature 2-form and $\epsilon_{ab}$ being the Levi-Civita symbol. Therefore, by inferring the parameter identifications [^2] Obviously, [G]=1[G]=-1 and $\Lambda]=1$.,
$G$ $=$ $\frac{1}{16\pi\mathrm{e}^{2}}\;,$
$\Lambda$ $=$ $\mathrm{e}^{2}\;,$ (3.3)
@@
This is a two-dimensional gravity action, in the first-order formalism, coupled to a scalar field naturally identified as the dilaton. The constants $G$ and $\Lambda$ are interpreted as the two-dimensional Newtonian and cosmological constants, respectively. Remarkably, since the electric charge is known, the numerical values of these gravitational constants can be computed through expressions (3.3). Indeed,
@@
The first term in the action (3.4) is the standard kinematic term for a real scalar field, up to a constant factor. The second and third terms are the typical Einstein-Hilbert (EH) and cosmological constant terms. The EH action, in two dimensions, is known to be topological, a property reflected by the fact that it reduces to a total derivative. In fact, in two dimensions, the EH term coincides with the GB term. See Section 5.
Continuing the analysis of the action (3.4), the next term is the dilaton-curvature coupling, which arises from the CS term. The last term is just a total derivative. This term, just like the EH one, could only contribute to the field equations at the boundary if this boundary is non-trivial.
Clearly, the action (3.4) is, up to boundary terms, manifestly gauge invariant under $SO(2)$ rotations in the tangent space,
@@
Thus, the action (3.4) is a genuine two-dimensional Euclidean gravity in the FOF, coupled to a scalar field. We point out that transformations (3.7) are directly obtained from (2.15), (3.1) and the identification $\alpha(x)=\epsilon_{ab}\alpha^{ab}(X)$, see [56, 57]. We also stress that action (3.4) exhibits diffeomorphism symmetry, which is implicit in the form notation.
In the next section, we will show that this action can be transmuted into JT-type gravities. For now, let us focus on the dynamics of the action (3.4). The field equation for $\Phi$ is just
@@
indicating that the scalar field carries some dynamics. It also suggests that the scalar field non-trivially influences the geometry of the manifold. However, the spin connection equation gives
@@
So, the dilaton solution is simply an arbitrary constant field configuration, possibly zero. Furthermore, combining equations (3.8) and (3.9) one easily finds that the manifold is locally flat:
@@
It remains to derive the equation of the zweibein. The only nontrivial term remaining in the action (3.4) that depends on the zweibein is the area term. Therefore, the final equation simply corresponds to the minimization of the area of the manifold (Plateau problem), just as two-dimensional Mardones-Zanelli gravity [35, 36]. The solution here is thus a flat space with minimal area. Of course, the specific manifold depends on the boundary conditions. Typical noncompact examples are planes, cylinders, and disks. In the case of compact manifolds, the flat torus is the standard solution. See Figure 1, in Section 5, for examples. In fact, we will return to the discussion of these surfaces in Section 5.
@@ 4.1 Hyperbolic solution
The first way to obtain JT gravity is by performing the curvature shift
@@
enforcing the usual JT term to appear directly in the gravity action (3.4). However, before performing this shift properly, let us define it a bit more rigorously. In fact, the shift (4.1) is actually generated by a shift in the spin connection, namely,
@@
meaning that the field $X$ must be a non-conservative field with a canonical dimension of [X]=1[X]=-1. For further reference, we name the condition (4.3) shift constraint. Thus, once the zweibein $e$ describing the manifold is chosen (note that there are infinite possibilities due to the $SO(2)$ gauge symmetry and diffeomorphisms), the field $X$ can be determined. Integrating equation (4.3) and applying Stokes’ theorem gives the global condition
@@
Another point of importance concerning equation (4.3) is that it is not gauge covariant unless $dX$ is covariant, which is actually not true. In fact, the shift (4.2) suggests that $X$ is a gauge invariant field. As a consequence, gauge symmetry is broken by the constraint (4.3); it may be seen, thus, as a gauge fixing for the model. In fact, together with the shift (4.2), it can be incorporated into the action (3.4) with a suitable 0-form Lagrange multiplier [^3] The Lagrange multiplier carries canonical dimension 3. $\lambda_{ab}$ by means of the addition of the action
@@
This term clearly breaks gauge symmetry, assuming that $\lambda$ is either gauge-invariant or covariant.
@@
Let us derive the field equations. The equation for the Lagrange multiplier naturally imposes the shift constraint (4.3). The equation of the spin connection remains unchanged, equation (3.9). Thus, $\Phi$ is again undetermined. It is constant, but not necessarily zero. However, nothing prevents us from choosing $\Phi=0$. The equation for $X$, provides (already employing (3.9))
@@
which corresponds to the JT solution for negative constant curvature in Euclidean spaces.
@@
This is actually quite an interesting result because a Lagrange multiplier functions as a force that enforces the system to satisfy the constraint. It turns out that the cosmological constant plays the role of that force. In essence, the cosmological constant is the force that bends the manifold.
The solution we found so far is just a trivial dilatonic vacuum and a manifold with negative constant curvature, i.e., a hyperbolic surface. In two-dimensional Euclidean spaces, such manifolds, if compact, are the connected $n$-holed tori [64, 65] - see Figure 3 and Section 5 for further discussion.
@@ 4.2 Spherical solution
Let us compute the field equations. Equation (3.9) remains valid, stating that $\Phi$ is a constant field again. The equation of $\lambda$ simply enforces the shift constraint (4.3). Again, the equation of the field $X$ gives that $d\lambda=0$, due to (3.9). Thus, $\lambda$ is also a constant field. The field equation for the dilaton, already employing the equations of $\omega$ and $\lambda$, gives the JT solution for a spherical manifold (constant positive curvature),
@@ 4.3 Extra remarks about the shift constraint
To conclude this Section, we remark that the shift constraint (4.3) can also be visualized as the imposition of a background. Therefore, the gauge fixing is actually a choice of an appropriate background. The only distinction between the hyperbolic and the spherical cases lies in the relative sign of the background. Essentially, the shifts (4.2) and (4.12), together with the constraint (4.3), impose the background to be hyperbolic or spherical. Moreover, since in two dimensions the EH term is topological, there is no room to derive Einstein-like equations that would provide dynamics around the background.
In fact, the shifts (4.2) and (4.12) are simply stating that $\sim\pm X$ is the spin connection background in each case. The shift constraint (4.3) states that $X$ must be the field that generates the specific (hyperbolic or spherical) background curvature. Moreover, it is known that the spin connection carries a specific term depending on the zweibein. Therefore, it is only natural that the shift constraint is written in terms of the zweibein. This is even evidenced by the fact that torsion is zero in our solutions.
@@
It is also interesting to notice that the JT term could emerge directly from the mapping by considering an electromagnetic background field. Right from the action (2.16), one could consider the shift $F\longrightarrow F\pm\Theta/2$. Or, equivalently, a shift on the electromagnetic field $a\longrightarrow a\pm M$ with $M$ being fixed by $dM=\Theta/2$. In this way, the target gravity action would already contain the usual JT term $\Phi(R\pm\Lambda^{2}/2)$. Hence, the shift we proposed is equivalent to the existence of a background electromagnetic field in the original MCS theory.
Another comment is that the shift constraint (4.3) induces a relation between the spin connection and the field $X$. This relation is actually generated by the cohomology of the nilpotent exterior derivative $d$, the well-known de Rham cohomology [66]. In fact, the constraint (4.3) together with the curvature equations (4.8) (negative sign) and (4.14) (positive sign) can be combined to deduce the on-shell relation
@@
Moreover, since $X$ is gauge invariant, $Y$ must transform like a connection in order to preserve the gauge structure of relation (4.16), namely
@@
Therefore, the gauge invariant field $X$ is essentially a difference between two connections, $X=\mp 2(\omega-Y)/\Lambda^{2}$, i.e., the difference between the spin connection and an arbitrary closed 1-form.
Condition (4.17) is actually a traditional cohomology problem whose solution is simply given by its non-trivial and trivial parts in de Rham cohomology, namely
@@
while $N$ is an arbitrary 0-form (and hence cannot be written as an exact quantity) with dimension $[N]=0$. The gauge transformations of $Z$ and $N$ can be inferred from (4.18) and (4.19). Clearly, $Z$ can be chosen to be a gauge connection while $N$ is set to be gauge invariant [^4] Alternatively, there is also the possibility to choose $Z$ to be a gauge invariant field while $N$ would transform in a non usual way by $\delta N=\alpha$, i.e., $N$ would translate under $SO(2)$ gauge transformations. Such transformation indicates that $N$ could be interpreted as a field of angular nature. In this case, the relation (4.19) could be seen as a gauge transformation of the field $Y$ while $N$ would be the gauge parameter.,
@@
The determination of the spin connection proceeds as follows. Once the zweibein describing the manifold is chosen, one can find any $X$ satisfying the shift constraint (4.3), a closed non-exact field $Z$, satisfying (4.20), and any gauge-invariant 0-form $N$. Then, using relations (4.16) and (4.19), the spin connection is completely determined. The same reasoning works to find $X$ by solving equation (4.8) (or (4.14)) for the spin connection. In essence, the spin connection and the field $X$ differ only by any non-trivial field $Z$ plus an exact field. Moreover, $Z$ and $N$ take no part in the solutions we found for JT gravity; see Equations (4.8) and (4.14). Alternatively, $X$ and $\omega$ can be deduced independently from equations (4.3) and (4.8) (or (4.14)), respectively. Then, $Z$ and $N$ can be inferred by comparing $X$ and $\omega$ via relation (4.16).
It is also important to stress that $Y$ must be non-vanishing because of the gauge transformations of $\omega$, $Y$ (see (3.7) and (4.17)), and the gauge-invariant field $X$. Otherwise, relation (4.16) would not be gauge covariant. Thus, if $Z\neq 0$, then $dN$ can be set to vanish. On the other hand, since $N$ is gauge-invariant, $Z$ cannot be set to zero. Therefore, from now on, we are allowed to fix $N=0$ with no loss of generality. As an alternative, if we do not care about gauge covariance (which is a valid choice since the presence of $X$ in the action already breaks gauge symmetry), we can set $Z=N=0$ and work with a gauge-invariant spin connection. However, we will not follow that approach.
@@
Fields $e^{a}$ $\omega^{ab}$ $\Phi$ $X$ $Y$ $Z$ NN $\lambda$ $G$ $\Lambda$ $\kappa$
Dimension 0 11 11 1-1 11 11 0 33 1-1 11 0
Rank 11 11 0 11 11 11 0 0 0 0 0
@@
The set of fundamental fields defined on $M^{2}$ is thus $\mathcal{F}=\{e,\omega,X,Z,\Phi,\lambda\}$, interlinked by the field equations
@@
$R^{ab}$ $=$ $\pm\frac{\Lambda^{2}}{2}e^{a}e^{b}\;,$
$\omega^{ab}$ $=$ $\pm\frac{\Lambda^{2}}{2}X^{ab}+Z^{ab}\;.$ (4.22)
The third equation in (4.22) establishes that, among $\omega$, $X$, and $Z$, we can neglect one of them in favor of the other two. The space of fundamental fields is then reduced, for instance, to $\mathcal{F}_{o}=\{e,X,Z,\Phi,\lambda\}$, where we have chosen to omit the spin connection.